OpenStudy (mary1999):

Decide weather the polygons are similar if so write the scale factor of triangle XAR triangle MNT XR=50, MN=27, NT=36

OpenStudy (mary1999):



OpenStudy (adrimit):

of course, they are similar (two angles match)

OpenStudy (adrimit):

the scale factor is \(\frac{\sqrt{27^2+36^2}}{50}\) \[27^2=0449 \\\;\;\;\;\;\;\;\;\;\;\;\;28 \\\;\;\;\;\;\;=0729\] \[36^2=0936 \\\;\;\;\;\;\;\;\;\;\;\;\;36 \\\;\;\;\;\;=1296\] \[1296+729=1025\] Therefore the scale factor is \(\frac{\sqrt{1025}}{50}\)